Respected sir & madum,
1] find the ratio of x:y:z,from the following equation :
x-2y+z=0 .....(1)
3x+2y-3z=0....(2)
2]solve the following simultaneous equatins by cross multiplication method
5x-3y=1 ..... (1)
x+2y=12......(2)
plz solve my difficulty as soon an possible...need simple explaination.
regards.
Binita.

Answer by stanbon(75887)

1] find the ratio of x:y:z,from the following equations :
x-2y+z=0 .....(1)
3x+2y-3z=0....(2)
------------
Find z:x by eliminating "y":
Add the two equations to get:
4x - 2z = 0
2z = 4x
z = 2x
z:x = 2:1
x:z = 1:2
---------------
Now find x:y by eliminating "z":
3x -6y + 3z = 0
3x + 2y - 3z = 0
---
Add to get:
6x -4y = 0
3x = 2y
x/y : 2/3
x:y = 2;3
----------------
x:y:z = ?
2:3:4
=============

2]solve the following simultaneous equations by cross multiplication method
5x-3y=1 ..... (1)
x+2y=12......(2)
---------
(5x-3y)/(x+2y) = 1/12
12(5x-3y) = x+2y
60x - 36y = x+2y
59x = 38y
x = (38/59)y
-----
Substitute to solve for "y"
(38/59)y + 2y = 12
(19/59)y + y = 6
(78/59)y = 6
y = 59*6/78
y = 59/13
--------------------
Substitute to solve for "x":
x = (38/59)(59/13)
x = 38/13
============================
Cheers,
Stan H.